OpenStudy (hoot_hoot):

How do you solve this: Determine the value of x so that the given points are collinear. (x + 1, 4), (2, 1), (2x + 15, -2) Please, help.

OpenStudy (eliesaab):

Equate slopes \[ \frac{1-4}{-x+2-1}=\frac{-2-1}{2 x+15-2} \]

OpenStudy (eliesaab):

Solve for x

OpenStudy (eliesaab):

You will get x =-4

OpenStudy (will.h):

we call the points collinear if they lay on the same straight line

OpenStudy (mathmale):

Think: How could you check your results?

OpenStudy (hoot_hoot):

Thank you. @eliesaab

OpenStudy (eliesaab):

YW